OpenStudy (agentjamesbond007):

What is the slope of the tangent line of y= (e^(-x))/(x+1) at x=1

OpenStudy (anonymous):

you need to find derivative y' = (x+1)*-e^(-x) - (e^(-1) -------------------- (x+1)^2 at x = 1 y' = 2*-e^-1 - (e^-1) -------------- 4 = e^(-1) / 4 = 0.092 slope of tangent = 0.092

OpenStudy (anonymous):

my maths is rubbish today - i made an error in last line y' = -3e^-1 / 4

OpenStudy (anonymous):

i get that = -1.1036

OpenStudy (anonymous):

divided by 4 which is -0.2759